Two 4.0 cm ×4.0 cm square aluminum electrodes, spaced 0.60 mm apart, are connected to a 400 V battery. What is the charge on the positive electrode?
Answer #1

Capacitance of the electrode , C =

= 8.854*10^-12 F/m, A = 4 cm*4 cm = 1.6*10^-3 m^2, d = 0.60 mm = 0.6*10^-3 m

therefore C = (8.854*10^-12 * 1.6*10^-3)/ 0.6*10^-3 F = 2.36*10^-11 F

Voltage, V = 400 v

we know charge in a capacitor, Q = C*V = 2.36*10^-11 * 400 C = 9.44 * 10^-9 C